# Constraining the Extragalactic Magnetic Field: Auger Data Meet UHECR Propagation Modeling

### v1

@@ 3.1 Probability density map
[…]
Figure 2: Probability density maps for the jetted AGN catalog in Galactic coordinates. The best fit was found at $\tilde{B}=2.9$ nG Mpc1/2 and $E_{\mathrm{th}}=41$ EeV (top left), and the secondary maximum at $\tilde{B}=4$ nG Mpc1/2 and $E_{\mathrm{th}}=58$ EeV (top right). The map (bottom right) for $\tilde{B}=1.4$ nG Mpc1/2 and $E_{\mathrm{th}}=60$ EeV is added as an additional example. The sky map in the bottom left panel is reproduced from Abreu et al. (2022) without modification for comparison purposes (published under the Creative Commons Attribution 4.0 International License). The supergalactic plane is shown as a gray line in all sky maps. These maps do not include the contribution from the isotropic background or the exposure of Auger, which are taken into account in the TS calculations.
@@
To observe the effect of deflection in the EGMF, we plot the probability density map at different magnetic-field strengths. In Fig. 1, we show the distribution of the best-fit sky map in Galactic coordinates for the SBG sources with $\tilde{B}=3.4$ nG Mpc1/2 and $E_{\mathrm{th}}=38$ EeV (top left). It demonstrates a localized clustering of the cosmic-ray flux, with visible deflection effects caused by the EGMF. This map can be compared with the best-fit model map of Auger (bottom left) above $40$ EeV, without modification (Abreu et al., 2022). This map of Auger does not account for distance-dependent deflection effects, but its distribution is similar to our maps as the dominant sources, M 82, NGC 4945, and NGC 253, are all at around 3.5 to 3.7 Mpc distance.
The top-right panel corresponds to a magnetic-field strength of $\tilde{B}=4$ nG Mpc1/2 at $E_{\mathrm{th}}=58$ EeV, which is a secondary local maximum in the parameter space. The clustering in this case is similar due to a stronger magnetic-field strength (leading to stronger deflections) and an increase in the energy threshold (decreasing the deflections), but is compared with a different data set, as the threshold energy is applied to the observed data as well. The bottom-right panel shows the reduction of deflections for a smaller magnetic-field strength, at $\tilde{B}=1.4$ nG Mpc1/2 and $E_{\mathrm{th}}=38$ EeV.
@@
In contrast to our approach, the Auger methodology assigned the same angular width of the Fisher distribution to all sources in the catalog, irrespective of the distance to the source. In particular, in our approach, the closest sources are also the brightest, forming a hot spot expected in the Auger field of view, particularly in the direction of the group of galaxies composed of NGC 4945, NGC 253, M 82, and Centaurus A (see Fig. 2). All sources depicted in the map are very luminous and are located at distances of $<4$ Mpc from Earth. The M 82 hotspot lies mostly outside the exposure region of the Auger, which makes it clearly observable by TA. The deflection caused by the magnetic field is minimal for small magnetic fields but increases as we raise the magnetic-field strength, as indicated in the panels.
Fig. 2 focuses on the jetted AGN catalog and similarly shows the distribution of the best-fit sky map, the secondary maximum, the Auger best-fit, and an additional example map, all in Galactic coordinates. The best-fit map is shown in the top-left panel, with $\tilde{B}=2.9$ nG Mpc1/2 and $E_{\mathrm{th}}=41$ EeV. This map highlights the concentration of flux near Centaurus A, the dominant source in this catalog, reflecting its strong contribution to the observed anisotropy. The top-right panel of Fig. 2 shows the secondary local maximum, with $\tilde{B}=4$ nG Mpc1/2 and $E_{\mathrm{th}}=58$ EeV. The clustering in these two cases is again similar due to a stronger magnetic field strength and an increase in the energy threshold. The bottom-right panel shows the probability density map for $\tilde{B}=1.4$ nG Mpc1/2 and $E_{\mathrm{th}}=60$ EeV to show the effect of the EGMF strength on the UHECR distribution.
The sky map on the bottom left panel of Fig. 2 is taken from Abreu et al. (2022) for comparison with our model maps. Auger’s map is significantly different from our model maps, as, in our case, the expected anisotropies in the arrival directions of jetted AGNs are dominated by Centaurus A alone, in contrast to the Auger analysis, which identifies significant contributions from Mrk 421 and NGC 1275 as well. The difference is because, in our maps, the UHECRs from distant sources are strongly deflected, while in Auger’s map, the spread around the source position is the same no matter what the distance to the source is. With distances for Cen A of about 3.7 Mpc, NGC 1275 of about 78 Mpc, and Mrk 421 of about 134 Mpc, this results in large spreads (increased by significant energy losses) in our models for NGC 1275 and Mrk 421.
@@ 3.2 Maximum likelihood ratio analysis
To assess the compatibility of the observed UHECR arrival directions with the model predictions, we apply an unbinned maximum likelihood method. This statistical approach estimates two free parameters: the signal fraction $\alpha$, representing the fraction of events attributed to the astrophysical sources in the catalog, and the effective magnetic field strength $\tilde{B}$, which governs the angular spread of UHECRs around their source directions due to magnetic deflections. The analysis is repeated across a range of energy thresholds, from 32 EeV to 80 EeV, in steps of 1 EeV, to explore how the fit quality and optimal parameters vary with energy. At higher energies, UHECR deflections are expected to be smaller, potentially leading to stronger correlations with astrophysical source distributions.
Fig. 3 shows the values of $\hat{\alpha}$ and $\hat{\tilde{B}}$ that maximize the TS for each energy threshold. Results are presented separately for the SBG and jetted AGN catalogs. The results for SBGs and AGNs from our model are depicted in blue, while the corresponding results from the Auger analysis are shown in red. This figure provides insight into how the signal strength and inferred magnetic deflection scale with energy, and which catalog yields a better fit to the observed UHECR distribution.
As illustrated in Fig. 3, the TS curve reveals a two-peaks structure as a function of the energy threshold (top panel). For SBGs (left), the best-fit peak is at a threshold energy of 39 EeV, while it is at 41 EeV for AGNs (right). The second peak at energies 58 EeV for both catalogs, which is consistent with the findings reported by Auger.
@@
In the left panel of Fig. 4, the results for the SBG catalog are shown. This plot highlights the region in the $(\alpha,\tilde{B})$ plane where the TS reaches its maximum. The dashed red and green contours correspond to the $68.3\%$ and $95.4\%$ C.L., respectively, derived from the TS distribution under the Wilks’ theorem approximation. The right panel presents the corresponding results for the jetted AGN catalog. As in the left panel, the best-fit parameters are marked with a cross, and the C.L. contours illustrate the uncertainties on the fit parameters. The best-fit parameters obtained above the energy threshold that maximizes the departure from isotropy are marked with a cross.
Figure 3: The test statistic (top), signal fraction (middle), and magnetic-field strength (bottom) when maximizing the deviation from isotropy for the covered range of threshold energies, for SBGs (left), and jetted AGNs (right). The results of our model (blue) are compared with the results from Abreu et al. (2022) (red) for the test statistic and corresponding value of the signal fraction. The vertical lines indicate the positions of the best fit and the secondary maximum. For the sake of readability, the uncertainties in the parameters, which are correlated above successive energy thresholds, are not displayed.
Figure 4: The test statistic values for SBGs (left) and jetted AGNs (right), as a function of the signal fraction and magnetic-field strength. The dashed green and red lines represent $68.3\%$ and $95.4\%$ C.L. contours (1 d.o.f.), respectively. The crosses indicate the best-fit positions.
@@
SBGs 38 $16^{+11}_{-9}$ $4^{+1}_{-2}$ 25.13
Jetted AGNs 41 $9^{+12}_{-3}$ $3^{+3}_{-1}$ 19.45
SBGs 58 $20^{+19}_{-7}$ $4^{+3}_{-2}$ 17.5
Jetted AGNs 58 $12^{+17}_{-6}$ $4^{+2}_{-2}$ 17.4
@@
From the two-dimensional likelihood scans presented in Fig 4, we also derive upper limits on the effective magnetic field strength $\tilde{B}$. By profiling the likelihood over the signal fraction $\alpha$ and identifying the boundary at which the test statistic decreases by 2.71 units from its maximum (corresponding to a 90% confidence level for one degree of freedom), we place constraints on the maximum allowable deflections consistent with the observed data. The resulting upper limits on the magnetic field strength are $\tilde{B}<4.4$ nG Mpc1/2 for the SBG catalog and $\tilde{B}<6.7$ nG Mpc1/2 for the jetted AGN catalog, both at the 90% confidence level. The resulting lower limits on the magnetic field strength are $\tilde{B}>2$ nG Mpc1/2 for both the SBG and jetted AGN catalogs, at the 90% confidence level.
The EGMG magnetic field strength limits obtained provide valuable insight into the extragalactic magnetic environment and the degree of deflection experienced by UHECRs en route to Earth.
@@ 4 Discussion
In this work, we have focused on the deflections that UHECRs incur during their propagation in the EGMF. Additional deflections of UHECRs, however, are expected in the GMF. The GMF consists of coherent and turbulent magnetic field components. Including the turbulent GMF would only have further increased the rms average deflection angle. Our results, therefore, provide robust upper limits on the EGMF strength. However, the lower limits on the EGMF are only applicable if the deflections in the GMF are negligible. The expected deflections in the GMF differ significantly between different GMF models. This can, for example, be seen from Unger & Farrar (2025) where expected deflections in the GMF are shown for eight different variations of the UF23 models (Unger & Farrar, 2024). Furthermore, should a magnetized Galactic halo is present around our Galaxy at larger scales than in the UF23 models, the expected deflections in the GMF increase (see e.g. Shaw et al. (2022)).
@@
Although we do not explicitly consider the deflection introduced by the GMF, our EGMF upper limit can be re-normalized to give correspondingly larger upper limits on the turbulent GMF. The level of this re-normalization is dependent on the propagation distance through the turbulent GMF (see Eqn. 3). As an example, for a Galactic halo of size $D=30$ kpc, and a coherence length of $l_{\rm coh}=1$ kpc, an EGMF upper limit of 4 nG obtained for sources with $D\approx 3.6$ Mpc would convert to an upper limit constraint on the turbulent GMF of $1.4~{}\mu$ [^2] Interestingly, for the case in which the UHECRs arriving to Earth are dominated by deflections in the Galactic halo, the deflection angles for particles with energies below 10 EeV would expect $\theta_{\rm rms}>100^{\circ}$, resulting in the observed skymap becoming dipolar-like, see Shaw et al. (2025).. This limit sits close to, and is compatible with, the present detection hints of the halo GMF (Kronberg et al., 2008; Bernet et al., 2008; Heesen et al., 2023).
@@
Our EGMF setup only includes turbulent EGMFs; no structured EGMFs were considered. Turbulent magnetic fields typically cause a spreading out of arrival directions around the original source position, but do not change the apparent source position (see, however, Dolgikh et al. (2025)). Structured magnetic fields typically do the opposite; they change the apparent source position but do not cause a spread of arrival directions around the apparent source position (see e.g. Shaw et al. (2022) for a discussion on these two effects). Our results, therefore, hold specifically for turbulent EGMFs with a maximum correlation length less than the typical source distances (3 to 4 Mpc for the nearest sources we consider). When comparing the overdensities in the UHECR arrival directions of the Auger data (Abreu et al., 2022) with the source positions by eye, a shift of the source positions does not seem to be necessary if the overdensities are dominated by the nearest sources in the catalogs within the field of view of the Auger (M 83, NGC 4945, Circinus, and NGC 253 for starburst galaxies, and Cen A for jetted AGN). However, for the jetted AGN case, the overdensity in the Galactic South is hard to explain with turbulent magnetic fields alone, which drives the higher significance for a deviation from isotropy for the starburst galaxies case over the jetted AGN case (although the preference for starburst galaxies over jetted AGN should not be considered statistically significant). Including the source distances and the related deflections in structured EGMFs carries further information which could be used for “tomographic” analyses if better statistics and UHECR composition measures are available in the future.
@@
For the deflection angle simulations, we fixed the source UHECR composition to nitrogen. Additionally, the source spectral index and maximum rigidity were set to values obtained from Batista et al. (2019). These choices have been found to provide a good fit to the spectrum and composition measured by Auger in the relevant energy range. Ehlert et al. (2023) have shown that the maximum rigidity cannot vary much from source to source to maintain a good fit to the Auger data. The UHECR composition, however, can be heavier than pure nitrogen, especially at the high energy end of the considered energy range (depending on the choice of hadronic interaction model). For the same EGMF, a heavier composition would increase the angular deflection, resulting in stronger upper limits on the EGMF being obtained. The choice for a nitrogen composition, therefore, is conservative for obtaining upper limits on the magnetic-field strength. A predominantly lighter composition than nitrogen for $E>32$ EeV is hard to reconcile with the Auger composition data. A subdominant light component might still exist, but would have to be produced by sources with a higher $R_{\mathrm{max}}$ than the dominant sources (assuming rigidity-dependent cutoffs). It is, therefore, not likely that this is produced by relatively nearby sources from one of the catalogs.
@@
The interactions of UHECRs with background EBL and CMB photon fields can play a significant role for the catalog sources. For nitrogen with an energy of $E\sim 30$ EeV, the typical energy-loss-length is $\gtrsim 200$ Mpc (see e.g. Hooper et al. (2007); Alves Batista et al. (2015)). At higher energies, this loss length exponentially decreases (Taylor et al., 2009), with the loss length for $E\sim 50~{}{\rm EeV}$ becoming $\sim 70$ Mpc. These energy losses significantly reduce the potential contribution from sources like Mrk 421 (at 133.7 Mpc distance in the catalog) and NGC 1275 (at 78.0 Mpc). In Auger’s analysis, these interactions are taken into account by applying energy-dependent and distance-dependent weights to the sources based on the expected attenuation. In our analysis, we use the same weights as computed by Auger for the normalization of each source to follow their analysis as closely as possible. However, in our simulations, the effect of energy losses is stronger for scenarios with stronger magnetic fields, as the effective propagation distance is increased when the deflections in magnetic fields become stronger, an effect of particular importance given the exponential decrease of the loss length noted above. In addition, cosmic rays with lower rigidities are more strongly deflected by magnetic fields, also effectively increasing their propagation distance. These effects introduce additional dependencies of the expected arrival directions on the magnetic-field strength and the energy threshold, which are not present in scenarios where a blurring based on the source distance is taken into account (e.g. Abdul Halim et al. (2024a)).
@@
In an earlier study, van Vliet et al. (2021) derived magnetic field constraints for SBGs using an independently developed foreground–background model that interpolates between cataloged sources (foreground) and an isotropic background beyond the listed sources. The local source density, $\rho_{0}$, was introduced—alongside the magnetic field strength, $\tilde{B}$—to parametrize the relative contribution of the isotropic background. Since $\rho_{0}$ can be directly compared to expectations from star-forming and spiral galaxies, robust lower limits on the magnetic field strength, $\tilde{B}\gtrsim 0.2,\mathrm{nG}$ Mpc1/2, could be derived for SBGs (see also Durrer & Neronov (2013) for an estimated lower limit on the EGMF strength from UHECRs). These limits are consistent with the constraints obtained in the present study. The origin of this lower limit lies in the requirement that either sufficiently strong magnetic fields must be present to produce the isotropic component of the map, or the source distribution itself must be intrinsically isotropic (i.e., isotropic background dominated).
@@
Other works stating upper limits on the strength of EGMFs based on UHECR data are, e.g., Bray & Scaife (2018); Bister & Farrar (2024); and Neronov et al. (2023). These papers state lower magnetic field strengths than obtained in our work, but are generally based on stronger assumptions (besides using different methods). Bray & Scaife (2018), for example, based their limits on a typical source distance of around 70 Mpc, while their results become comparable to our upper limit for a scenario where they assumed a typical source distance of 10 Mpc. Rather than using the Auger catalog, which motivated the original correlation, Bister & Farrar (2024) assumed UHECR sources following the local matter density. This assumption leads to an upper limit on the EGMF for specific source densities. In addition, they included GMF deflections in a specific GMF model, instead of neglecting the GMF deflections, to obtain their upper limits on the EGMF. Neronov et al. (2023) obtained their upper limits on the EGMF assuming that the TA excess (Abbasi et al., 2021) is produced by the Perseus-Pisces supercluster at a distance of about 70 Mpc. Unfortunately, Auger has not been able to confirm the existence of an excess in the same direction, despite having a comparable integrated exposure in that region (Abdul Halim et al., 2025a).
Note that our derived EGMF constraints relate to the EGMF between the (studied) local galaxies and the Milky Way, not to EGMF voids in general. The distances to the dominant local sources (M 82, NGC 4945, NGC 253, and Cen A) are typically $<4$ Mpc. The larger typical source distances assumed in Bray & Scaife (2018); Bister & Farrar (2024); and Neronov et al. (2023) correspond more closely to limits on EGMFs in voids.
@@
We have also demonstrated that while for SBG the same sources as for the Auger analysis dominate the correlations, observable correlations with jetted AGNs other than Cen A in the corresponding source catalog are eliminated by the EGMF once the source distances and corresponding deflections are included (similar to what was found in Abdul Halim et al. (2024a)). These particular sources could help to discriminate between potential Galactic and extragalactic magnetic fields or isolate the regions with the strongest deflections in the future.
The future of UHECR research looks promising with the advancements brought by AugerPrime (Aab et al., 2016) and the Global Cosmic Ray Observatory (GCOS) initiative (Hörandel et al., 2021). AugerPrime, an ongoing upgrade to the Pierre Auger Observatory, will significantly enhance the precision of UHECR energy and composition measurements, potentially allowing the composition to be measured on a shower-by-shower basis. Meanwhile, the GCOS project aims to establish a global network of observatories, including AugerPrime. By combining data from multiple observatories, GCOS will help identify correlations between UHECRs and astrophysical sources, such as active galactic nuclei and gamma-ray bursts. Together, AugerPrime and GCOS will push the boundaries of cosmic ray research, enabling scientists to investigate the origins, propagation, and anisotropies of UHECRs, and make strides in answering some of the most fundamental questions in astrophysics.

### v2

@@ 3.1 Probability density map
Figure 1: Probability density map for the SBG catalog in Galactic coordinates. The best fit was found at $\tilde{B}=\;3.4$ nG Mpc1/2 and $E_{\mathrm{th}}=38$ EeV (top left), and the secondary maximum at $\tilde{B}=\;4$ nG Mpc1/2 and $E_{\mathrm{th}}=58$ EeV (top right). The map (bottom left) for $\tilde{B}=0.5$ nG Mpc1/2 and $E_{\mathrm{th}}=38$ EeV is added as an additional example to show the effect of a reduced magnetic-field strength. The sky map in the bottom-right panel is reproduced from Abreu_2022 without modification for comparison purposes (published under the Creative Commons Attribution 4.0 International License). The supergalactic plane is shown as a gray line in all sky maps. Our model maps are depicted in blue to distinguish between the Auger map and our models. These maps do not include the contribution from the isotropic background or the exposure of Auger, which are taken into account in the TS calculations.
Figure 2: Probability density maps for the jetted AGN catalog in Galactic coordinates. The best fit was found at $\tilde{B}=2.9$ nG Mpc1/2 and $E_{\mathrm{th}}=41$ EeV (top left), and the secondary maximum at $\tilde{B}=4$ nG Mpc1/2 and $E_{\mathrm{th}}=58$ EeV (top right). The map (bottom left) for $\tilde{B}=0.5$ nG Mpc1/2 and $E_{\mathrm{th}}=60$ EeV is added as an additional example. The sky map in the bottom right panel is reproduced from Abreu_2022 without modification for comparison purposes (published under the Creative Commons Attribution 4.0 International License). The supergalactic plane is shown as a gray line in all sky maps. These maps do not include the contribution from the isotropic background or the exposure of Auger, which are taken into account in the TS calculations.
@@
To observe the effect of deflection in the EGMF, we plot the probability density map at different magnetic-field strengths. In Fig. 1, we show the distribution of the best-fit sky map in Galactic coordinates for the SBG sources with $\tilde{B}=3.4$ nG Mpc1/2 and $E_{\mathrm{th}}=38$ EeV (top left). It demonstrates a localized clustering of the cosmic-ray flux, with visible deflection effects caused by the EGMF. This map can be compared with the best-fit model map of Auger (bottom right) above $40$ EeV, without modification (Abreu_2022). This map of Auger does not account for distance-dependent deflection effects, but its distribution is similar to our maps as the dominant sources, M 82, NGC 4945, and NGC 253, are all at around 3.5 to 3.7 Mpc distance.
The top-right panel corresponds to a magnetic-field strength of $\tilde{B}=4$ nG Mpc1/2 at $E_{\mathrm{th}}=58$ EeV, which is a secondary local maximum in the parameter space. The clustering in this case is similar due to a stronger magnetic-field strength (leading to stronger deflections) and an increase in the energy threshold (decreasing the deflections), but is compared with a different data set, as the threshold energy is applied to the observed data as well. The bottom-left panel shows the reduction of deflections for a smaller magnetic-field strength, at $\tilde{B}=0.5$ nG Mpc1/2 and $E_{\mathrm{th}}=38$ EeV.
@@
In contrast to our approach, the Auger methodology assigned the same angular width of the Fisher distribution to all sources in the catalog, irrespective of the distance to the source. In particular, in our approach, the closest sources are also the brightest, forming a hot spot expected in the Auger field of view, particularly in the direction of the group of galaxies composed of NGC 4945, NGC 253, M 82, and Centaurus A (see Fig. 2). All sources depicted in the map are very luminous and are located at distances of <4<4 Mpc from Earth. The M 82 hotspot lies mostly outside the exposure region of the Auger, which makes it clearly observable by TA. The deflection caused by the magnetic field is minimal for small magnetic fields but increases as we raise the magnetic-field strength, as indicated in the panels.
Fig. 2 focuses on the jetted AGN catalog and similarly shows the distribution of the best-fit sky map, the secondary maximum, the Auger best-fit, and an additional example map, all in Galactic coordinates. The best-fit map is shown in the top-left panel, with $\tilde{B}=2.9$ nG Mpc1/2 and $E_{\mathrm{th}}=41$ EeV. This map highlights the concentration of flux near Centaurus A, the dominant source in this catalog, reflecting its strong contribution to the observed anisotropy. The top-right panel of Fig. 2 shows the secondary local maximum, with $\tilde{B}=4$ nG Mpc1/2 and $E_{\mathrm{th}}=58$ EeV. The clustering in these two cases is again similar due to a stronger magnetic field strength and an increase in the energy threshold. The bottom-left panel shows the probability density map for $\tilde{B}=0.5$ nG Mpc1/2 and $E_{\mathrm{th}}=60$ EeV to show the effect of the EGMF strength on the UHECR distribution.
The sky map on the bottom right panel of Fig. 2 is taken from Abreu_2022 for comparison with our model maps. Auger’s map is significantly different from our model maps, as, in our case, the expected anisotropies in the arrival directions of jetted AGNs are dominated by Centaurus A alone, in contrast to the Auger analysis, which identifies significant contributions from Mrk 421 and NGC 1275 as well. The difference is because, in our maps, the UHECRs from distant sources are strongly deflected, while in Auger’s map, the spread around the source position is the same no matter what the distance to the source is. With distances for Cen A of about 3.7 Mpc, NGC 1275 of about 78 Mpc, and Mrk 421 of about 134 Mpc, this results in large spreads (increased by significant energy losses) in our models for NGC 1275 and Mrk 421.
@@ 3.2 Maximum likelihood ratio analysis
Using the unbinned maximum likelihood method described in Section 2.3, we estimate the best-fit signal fraction $\alpha$ and effective magnetic field strength $\tilde{B}$ for each energy threshold. This allows us to evaluate how the correlation between UHECR arrival directions and catalog sources evolves with energy.
Fig. 3 shows the values of $\alpha$ and $\tilde{B}$ that maximize the TS for each energy threshold. Results are presented separately for the SBG and jetted AGN catalogs. The results for SBGs and AGNs from our model are depicted in blue, while the corresponding results from the Auger analysis are shown in red. This figure provides insight into how the signal fraction and inferred magnetic deflection scale with energy, and which catalog yields a better fit to the observed UHECR distribution.
As illustrated in Fig. 3, the TS curve reveals a two-peak structure as a function of the energy threshold (top panel). For SBGs (left), the best-fit peak is at a threshold energy of 39 EeV, while for AGNs (right) it is at 41 EeV. The second peak, at energies of 58 EeV for both catalogs, is consistent with the findings reported by Auger.
[…]
@@
[…]
Figure 4: The test statistic values for SBGs (left) and jetted AGNs (right) as a function of the signal fraction and magnetic-field strength, evaluated at a threshold energy of 38 EeV for SBGs and 41 EeV for jetted AGNs. The dashed green and red lines represent $68.3\%$ and $95.4\%$ C.L. contours (1 d.o.f.), respectively. The crosses indicate the best-fit positions.
Catalog / Maximum $E_{\mathrm{th}}$ [EeV] Signal fraction ($\alpha$) $\%$] $\tilde{B}$ [nG Mpc1/2] TS
SBGs
Global maximum 38 $16^{+11}_{-9}$ $4^{+1}_{-2}$ 25.13
Secondary maximum 58 $19^{+10}_{-11}$ $4^{+3}_{-2}$ 17.50
Jetted AGNs
Global maximum 41 $9^{+12}_{-3}$ $3^{+3}_{-1}$ 19.45
Secondary maximum 58 $11^{+11}_{-5}$ $4^{+2}_{-2}$ 17.40
@@
From the two-dimensional likelihood scans presented in Fig 4, we also derive upper limits on the effective magnetic field strength $\tilde{B}$. By profiling the likelihood over the signal fraction $\alpha$ and identifying the boundary at which the test statistic decreases by 2.71 units from its maximum (corresponding to a 90% confidence level for one degree of freedom), we place constraints on the maximum allowable deflections consistent with the observed data. The resulting upper limits on the magnetic field strength are $\tilde{B}<4.4$ nG Mpc1/2 for the SBG catalog and $\tilde{B}<6.7$ nG Mpc1/2 for the jetted AGN catalog, both at the 90% confidence level. The resulting lower limits on the magnetic field strength are $\tilde{B}>2$ nG Mpc1/2 for both the SBG and jetted AGN catalogs, at the 90% confidence level. The EGMF strength limits obtained provide valuable insight into the extragalactic magnetic environment and the degree of deflection experienced by UHECRs en route to Earth.
@@ 4 Discussion
In this work, we have focused on the deflections that UHECRs incur during their propagation in the EGMF. Additional deflections of UHECRs, however, are expected in the GMF. The GMF consists of coherent and turbulent magnetic field components. Including the turbulent GMF would only have further increased the rms average deflection angle. Our results, therefore, provide robust upper limits on the EGMF strength (see e.g. Bray:2018). However, the lower limits on the EGMF are only applicable if the deflections in the GMF are negligible. The expected deflections in the GMF differ significantly between different GMF models. This can, for example, be seen from Unger:2025xkk where expected deflections in the GMF are shown for eight different variations of the UF23 models (unger2024). Furthermore, should a magnetized Galactic halo is present around our Galaxy at larger scales than in the UF23 models, the expected deflections in the GMF increase  (see e.g. Shaw:2022lqd).
@@
Although we do not explicitly consider the deflection introduced by the GMF, our EGMF upper limit can be re-normalized to give correspondingly larger upper limits on the turbulent GMF. The level of this re-normalization is dependent on the propagation distance through the turbulent GMF (see Eqn. 3). As an example, for a Galactic halo of size $D=30$ kpc, and a coherence length of $l_{\rm coh}=1$ kpc, an EGMF upper limit of 4 nG obtained for sources with $D\approx 3.6$ Mpc would convert to an upper limit constraint on the turbulent GMF of $1.4~\mu$ [^3] Interestingly, for the case in which the UHECRs arriving to Earth are dominated by deflections in the Galactic halo, the deflection angles for particles with energies below 10 EeV would expect $\theta_{\rm rms}>100^{\circ}$, resulting in the observed skymap becoming dipolar-like, see Shaw:2025ykm.. This limit sits close to, and is compatible with, the present detection hints of the halo GMF (Kronberg:2007dy; Bernet:2008qp; Heesen_2023).
Our EGMF setup only includes turbulent EGMFs; no structured EGMFs were considered. Turbulent magnetic fields typically cause a spreading out of arrival directions around the original source position, provided the distance traveled is significantly larger than the coherence length (Dolgikh:2025bac).
@@
Structured magnetic fields typically do the opposite; they change the apparent source position but do not cause a spread of arrival directions around the apparent source position (see e.g. Shaw:2022lqd for a discussion on these two effects). Our results, therefore, hold specifically for turbulent EGMFs with a maximum correlation length less than the typical source distances (3 to 4 Mpc for the nearest sources we consider). When comparing the overdensities in the UHECR arrival directions of the Auger data (Abreu_2022) with the source positions by eye, a shift of the source positions does not seem to be necessary if the overdensities are dominated by the nearest sources in the catalogs within the field of view of the Auger (M 83, NGC 4945, Circinus, and NGC 253 for starburst galaxies, and Cen A for jetted AGN). However, for the jetted AGN case, the overdensity in the Galactic South is hard to explain with turbulent magnetic fields alone, which drives the higher significance for a deviation from isotropy for the starburst galaxies case over the jetted AGN case (although the preference for starburst galaxies over jetted AGN should not be considered statistically significant). Including the source distances and the related deflections in structured EGMFs carries further information which could be used for “tomographic” analyses if better statistics and UHECR composition measures are available in the future.
@@
For the deflection angle simulations, we fixed the source UHECR composition to nitrogen. Additionally, the source spectral index and maximum rigidity were set to values obtained from Batista_2019. These choices have been found to provide a good fit to the spectrum and composition measured by Auger in the relevant energy range. Ehlert:2022jmy have shown that the maximum rigidity cannot vary much from source to source to maintain a good fit to the Auger data. The UHECR composition, however, can be heavier than pure nitrogen, especially at the high energy end of the considered energy range (depending on the choice of hadronic interaction model). For the same EGMF, a heavier composition would increase the angular deflection, resulting in stronger upper limits on the EGMF being obtained. The choice for a nitrogen composition, therefore, is conservative for obtaining upper limits on the magnetic-field strength. A predominantly lighter composition than nitrogen for $E>32$ EeV is hard to reconcile with the Auger composition data. A subdominant light component might still exist, but would have to be produced by sources with a higher $R_{\mathrm{max}}$ than the dominant sources (assuming rigidity-dependent cutoffs). It is, therefore, not likely that this is produced by relatively nearby sources from one of the catalogs.
@@
The interactions of UHECRs with background EBL and CMB photon fields can play a significant role for the catalog sources. For nitrogen with an energy of $E\sim 30$ EeV, the typical energy-loss-length is $\gtrsim 200$ Mpc (see e.g. Hooper:2006tn; AlvesBatista:2015jem). At higher energies, this loss length exponentially decreases (Taylor:2009iw), with the loss length for $E\sim 50~{\rm EeV}$ becoming $\sim 70$ Mpc. These energy losses significantly reduce the potential contribution from sources like Mrk 421 (at 133.7 Mpc distance in the catalog) and NGC 1275 (at 78.0 Mpc). In Auger’s analysis, these interactions are taken into account by applying energy-dependent and distance-dependent weights to the sources based on the expected attenuation. In our analysis, we use the same weights as computed by Auger for the normalization of each source to follow their analysis as closely as possible. However, in our simulations, the effect of energy losses is stronger for scenarios with stronger magnetic fields, as the effective propagation distance is increased when the deflections in magnetic fields become stronger, an effect of particular importance given the exponential decrease of the loss length noted above. In addition, cosmic rays with lower rigidities are more strongly deflected by magnetic fields, also effectively increasing their propagation distance. These effects introduce additional dependencies of the expected arrival directions on the magnetic-field strength and the energy threshold, which are not present in scenarios where a blurring based on the source distance is taken into account (e.g. PierreAuger:2023htc).
@@
In an earlier study, van_Vliet_2021 derived magnetic field constraints for SBGs using an independently developed foreground–background model that interpolates between cataloged sources (foreground) and an isotropic background beyond the listed sources. The local source density, $\rho_{0}$, was introduced—alongside the magnetic field strength, $\tilde{B}$—to parametrize the relative contribution of the isotropic background. Since $\rho_{0}$ can be directly compared to expectations from star-forming and spiral galaxies, robust lower limits on the magnetic field strength, $\tilde{B}\gtrsim 0.2,\mathrm{nG}$ Mpc1/2, could be derived for SBGs  (see also Durrer:2013pga) for an estimated lower limit on the EGMF strength from UHECRs). These limits are consistent with the constraints obtained in the present study. The origin of this lower limit lies in the requirement that either sufficiently strong magnetic fields must be present to produce the isotropic component of the map, or the source distribution itself must be intrinsically isotropic (i.e., isotropic background dominated).
@@
Other works stating upper limits on the strength of EGMFs based on UHECR data are, e.g., Bray_2018; Bister:2023icg; and Neronov:2021xua. These papers state lower magnetic field strengths than obtained in our work, but are generally based on stronger assumptions (besides using different methods). Bray_2018, for example, based their limits on a typical source distance of around 70 Mpc, while their results become comparable to our upper limit for a scenario where they assumed a typical source distance of 10 Mpc. Rather than using the Auger catalog, which motivated the original correlation, Bister:2023icg assumed UHECR sources following the local matter density. This assumption leads to an upper limit on the EGMF for specific source densities. In addition, they included GMF deflections in a specific GMF model, instead of neglecting the GMF deflections, to obtain their upper limits on the EGMF.
Neronov:2021xua obtained their upper limits on the EGMF assuming that the TA excess (TelescopeArray:2021dfb) is produced by the Perseus-Pisces supercluster at a distance of about 70 Mpc. Unfortunately, Auger has not been able to confirm the existence of an excess in the same direction, despite having a comparable integrated exposure in that region (PierreAuger:2024hrj). We note that, even if the TA excess is real, there may exist sources much closer than 70 Mpc, such as NGC 891 (Anchordoqui:2022pzo), or even M31 (Plotko:2022urd), which could contribute significantly to the observed flux.
Our derived EGMF constraints relate to the EGMF between the (studied) local galaxies and the Milky Way. These constraints are not applicable to EGMF voids in general. The distances to the dominant local sources (M 82, NGC 4945, NGC 253, and Cen A) are typically <4<4 Mpc. The larger typical source distances assumed in Bray_2018; Bister:2023icg; and Neronov:2021xua correspond more closely to limits on EGMFs in voids.
@@
We have also demonstrated that while for SBG the same sources as for the Auger analysis dominate the correlations, observable correlations with jetted AGNs other than Cen A in the corresponding source catalog are eliminated by the EGMF once the source distances and corresponding deflections are included (similar to what was found in PierreAuger:2023htc). These particular sources could help to discriminate between potential Galactic and extragalactic magnetic fields or isolate the regions with the strongest deflections in the future.
The future of UHECR research looks promising with the advancements brought by AugerPrime (PierreAuger:2016qzd) and the Global Cosmic Ray Observatory (GCOS) initiative (GCOS:2021exh). AugerPrime, an ongoing upgrade to the Pierre Auger Observatory, will significantly enhance the precision of UHECR energy and composition measurements, potentially allowing the composition to be measured on a shower-by-shower basis. Meanwhile, the GCOS project aims to establish a global network of observatories, including AugerPrime. By combining data from multiple observatories, GCOS will help identify correlations between UHECRs and astrophysical sources, such as active galactic nuclei and gamma-ray bursts. Together, AugerPrime and GCOS will push the boundaries of cosmic ray research, enabling scientists to investigate the origins, propagation, and anisotropies of UHECRs, and make strides in answering some of the most fundamental questions in astrophysics.